Diego Motors is a small car dealership. On average, it sells a car for
$25,000, which it purchases from the manufacturer for $22,000. Each month, Diego Motors pays $50,000 in
rent and utilities and $60,000 for salespeople’s salaries. In addition to their salaries, salespeople are paid a
commission of $500 for each car they sell. Diego Motors also spends $10,000 each month for local advertisements.
Its tax rate is 40%.
1. How many cars must Diego Motors sell each month to break even?
2. Diego Motors has a target monthly net income of $54,000. What is its target monthly operating income?
How many cars must be sold each month to reach the target monthly net income of $54,000?

Answer by stanbon(75887)

Diego Motors is a small car dealership. On average, it sells a car for
$25,000, which it purchases from the manufacturer for $22,000.
----------------
Let # of cars sold per month be "x"
----------
Net car income = (25000-22000)x dollars
---------------------
Each month, Diego Motors pays $50,000 in rent and utilities and $60,000 for salespeople’s salaries.
----------
Operations overhead = 50000+60000= 110,000 dollars
---------------
In addition to their salaries, salespeople are paid a
commission of $500 for each car they sell.
-------------
Commission costs = 500x dollars
------------------
Diego Motors also spends $10,000 each month for local advertisements.
----------
Total operating expenses: 10,000+500x+110,000 = 120000+500x dollars
Gross Income = 3000x
--------------------------
Gross profit 3000x-(120000+500x)= 2500x-120000
----------------
Its tax rate is 40%.
Net profit after taxes: 0.60(2500x-120000) = 1500x-72000 dollars
-------------------------
1. How many cars must Diego Motors sell each month to break even?
1500x-72000=0
x = 48 cars
---------------------------
2. Diego Motors has a target monthly net income of $54,000. What is its target monthly operating income?
1500x-72000 = 54000
1500x = 126000
x = 84 (number of cars it needs to sell to clear $54000)
You'll have to figure out its "target monthly operating income".
====================
Cheers,
Stan H.